22. Hen Hao Professor Mullen, Inc applies overhead on the basis of machine hours. Given the following data, compute overhead skip 23,24,26
Answer #1

Okay. I'm solving 22 and 25

22. Overhead rate = 1,200,000/400,000 = 3 per mh

Applied overhead = 3*380,000 = 1,140,000

Underapplied overhead = 1,155,000 - 1,140,000 = 15,000

Option C

25. Option B